# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 3.6 Rigorous PP analysis
Let the wave speeds in the HLL flux satisfy the condition (57). If the polynomial vectors $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ are locally divergence-free and satisfy the PP limiting condition (54), then the scheme (55) preserves $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ under the CFL-type condition
$\Delta t_{n}\leq c_{0}\min\left\{\frac{\Delta x\omega_{1}^{+}}{\alpha_{1,i+\frac{1}{2}}^{\mu}},\frac{\Delta x\omega_{1}^{-}}{\alpha_{1,i-\frac{1}{2}}^{\mu}},\frac{\Delta y\omega_{2}^{+}}{\alpha_{2,j+\frac{1}{2}}^{\mu}},\frac{\Delta y\omega_{2}^{-}}{\alpha_{2,j-\frac{1}{2}}^{\mu}},\right\}\quad\forall I_{ij}\in\Omega,1\leq\mu\leq q,$ (61)
where $c_{0}$ is the maximum allowable CFL number, and
$\alpha_{1,i+\frac{1}{2}}^{\mu}=\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}-\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}+\frac{\left|\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}},\quad\alpha_{1,i-\frac{1}{2}}^{\mu}=\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}+\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}},$
$\alpha_{2,j+\frac{1}{2}}^{\mu}=\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}-\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}+\frac{\left|\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}},\quad\alpha_{2,j-\frac{1}{2}}^{\mu}=\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}+\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}.$
We now present the proof of Theorem 1.
Proof According to the identities (58)-(59) in Corollary 2, the scheme (55) can be written as
$\bar{\boldsymbol{U}}_{ij}^{n+1}=\bar{\boldsymbol{U}}_{ij}^{n}-$ $\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}\boldsymbol{H}_{1,i+\frac{1}{2}}^{\mu}+\boldsymbol{F}_{1}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)\right.$ (62)
@@
$\left.+\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)-\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\right]$
@@
$\left.+\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)-\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\right].$
The scheme (62) above can be reformulated as
$\bar{\boldsymbol{U}}_{ij}^{n+1}=\bar{\boldsymbol{U}}_{ij}^{n}+\boldsymbol{\Xi}_{1}+\boldsymbol{\Xi}_{2}+\boldsymbol{\Xi}_{3}+\boldsymbol{\Xi}_{4},$ (63)
@@
$\boldsymbol{\Xi}_{4}:=$ $-\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)-\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\right)$
$-\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)-\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\right).$
We define
$\bar{\boldsymbol{U}}_{ij}^{(\mu)}=\frac{1}{\Delta y\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)+\Delta x\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)}\left[\Delta y\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}-\boldsymbol{F}_{1}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)\right)\right.$
$+\Delta y\left(\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}+\boldsymbol{F}_{1}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\right)+\Delta x\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}-\boldsymbol{F}_{2}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)\right)$
$+\left.\Delta x\left(\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}+\boldsymbol{F}_{2}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\right)\right],\quad\mu=1,2,\cdots,q,$
@@
$\boldsymbol{\Xi}_{2}=\frac{\Delta t_{n}}{\Delta x\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\Delta y\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)+\Delta x\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)\right)\bar{\boldsymbol{U}}_{ij}^{(\mu)}.$
Now, we first show $\bar{\rho}_{ij}^{n+1}>0$. Thanks to Corollary 1, we have, for all $1\leq\mu\leq q$, $\bar{\boldsymbol{U}}_{ij}^{(\mu)}\in\mathcal{G}_{\rho}$, and $\Delta y\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)+\Delta x\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)>0$. Therefore, $\boldsymbol{\Xi}_{2}\in\mathcal{G}_{\rho}$. According to Corollary 2, we have $\boldsymbol{H}_{1,i+\frac{1}{2}}^{\mu}$, $\boldsymbol{H}_{1,i-\frac{1}{2}}^{\mu}$, $\boldsymbol{H}_{2,j+\frac{1}{2}}^{\mu}$ and $\boldsymbol{H}_{2,j-\frac{1}{2}}^{\mu}$ all belong to the $\mathcal{G}_{\rho}$, which means $\boldsymbol{\Xi}_{3}\in\mathcal{G}_{\rho}$. Since the first component of $\boldsymbol{\Xi}_{4}$ is zero, we deduce from (63) that
@@
where we have used the convex decomposition (47) in the above equality, the PP limiting condition (54) in the first inequality, and the CFL-type condition (61) in the last inequality.
we then prove that $\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{{\left|\boldsymbol{B}^{*}\right|}^{2}}{2}>0$ for any $\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}$. It follows from (63) that
$\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}=\Pi_{0}+\Pi_{1}+\Pi_{2}+\Pi_{3}+\Pi_{4},$ (64)
@@
$\Pi_{0}:=\bar{\boldsymbol{U}}_{ij}^{n}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2},$
@@
$\Pi_{1}:=$ $\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}\right)\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right.$
$\left.-\left(\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right]$
@@
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}\right)\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right.$
$\left.-\left(\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right],$
@@
$\Pi_{2}:=\frac{\Delta t_{n}}{\Delta x\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\Delta y\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)+\Delta x\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)\right)\left(\bar{\boldsymbol{U}}_{ij}^{(\mu)}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right),$
$\Pi_{3}:=$ $-\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}\left(\boldsymbol{H}_{1,i+\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)-\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}\left(\boldsymbol{H}_{1,i-\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right]$
$-\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}\left(\boldsymbol{H}_{2,j+\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)-\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}\left(\boldsymbol{H}_{2,j-\frac{1}{2}}^{\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right],$
$\Pi_{4}:=$ $-\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i+\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)\cdot\boldsymbol{n}^{*}-\mathscr{B}_{1,i-\frac{1}{2}}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\cdot\boldsymbol{n}^{*}\right)$
$-\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j+\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)\cdot\boldsymbol{n}^{*}-\mathscr{B}_{2,j-\frac{1}{2}}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\cdot\boldsymbol{n}^{*}\right).$
We now estimate the lower bounds of $\Pi_{0}$, $\Pi_{2}$, $\Pi_{3}$ and $\Pi_{4}$, respectively. Based on the the convex decomposition (47), we have
$\Pi_{0}=$ $\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\omega_{1}^{+}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)+\omega_{1}^{-}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)+\omega_{2}^{+}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right.$ (65)
$\left.+\omega_{2}^{-}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right]+\sum_{s=1}^{S}\omega_{s}\left(\widetilde{\boldsymbol{U}}_{\sigma}^{n}\left(x_{ij}^{(s)},y_{ij}^{(s)}\right)\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right),$
$\geq\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\omega_{1}^{+}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)+\omega_{1}^{-}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right.$
$\quad\left.+\omega_{2}^{+}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)+\omega_{2}^{-}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left|\boldsymbol{B}^{*}\right|^{2}}{2}\right)\right],$
where the inequality follows from the PP limiting condition (54) and the Lemma 1. According to Corollary 1, $\forall\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}$, we have

### v2

@@ 3.6 Rigorous PP analysis
Let $\left\{\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)\right\}$ denote the OEDG solution polynomial vector in cell $I_{ij}$ after the OE and PP limiting procedures, so that $\widetilde{\boldsymbol{B}}_{ij}^{n}\left(\boldsymbol{x}\right)$ is LDF and $\widetilde{\boldsymbol{U}}_{ij}^{n}\left(\boldsymbol{x}\right)$ satisfies condition (58). If the wave speeds in the HLL flux satisfy the condition (62), then the scheme (59) preserves $\bar{\boldsymbol{U}}_{ij}^{n+1}\in\mathcal{G}$ under the CFL-type condition
$\Delta t_{n}\leq\min\left\{\frac{\Delta x\omega_{1}^{+}}{\alpha_{1,i+\frac{1}{2}}^{\mu}},\frac{\Delta x\omega_{1}^{-}}{\alpha_{1,i-\frac{1}{2}}^{\mu}},\frac{\Delta y\omega_{2}^{+}}{\alpha_{2,j+\frac{1}{2}}^{\mu}},\frac{\Delta y\omega_{2}^{-}}{\alpha_{2,j-\frac{1}{2}}^{\mu}}\right\}\quad\forall i,j,1\leq\mu\leq q,$ (66)
where
$\alpha_{1,i+\frac{1}{2}}^{\mu}=\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}-\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}+\frac{\left|\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i+\frac{1}{2},j}^{-,\mu}}},\quad\alpha_{1,i-\frac{1}{2}}^{\mu}=\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}+\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\right|}{\sqrt{\rho_{i-\frac{1}{2},j}^{+,\mu}}},$
$\alpha_{2,j+\frac{1}{2}}^{\mu}=\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}-\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}+\frac{\left|\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j+\frac{1}{2}}^{\mu,-}}},\quad\alpha_{2,j-\frac{1}{2}}^{\mu}=\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}+\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}+\frac{\left|\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\right|}{\sqrt{\rho_{i,j-\frac{1}{2}}^{\mu,+}}}.$
Proof According to the identities (63)–(64) in Lemma 4, the scheme (59) can be rewritten as
$\bar{\boldsymbol{U}}_{ij}^{n+1}=\bar{\boldsymbol{U}}_{ij}^{n}-$ $\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}\boldsymbol{H}_{1,i+\frac{1}{2}}^{\mu}+\boldsymbol{F}_{1}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)-\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)\right.$
@@
$\left.+\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)+\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\right]$
@@
$\left.+\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)+\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\right].$
It can be further split into
$\bar{\boldsymbol{U}}_{ij}^{n+1}=\bar{\boldsymbol{U}}_{ij}^{n}+\boldsymbol{\Xi}_{1}+\boldsymbol{\Xi}_{2}+\boldsymbol{\Xi}_{3}+\boldsymbol{\Xi}_{4}$ (67)
@@
$\boldsymbol{\Xi}_{4}:=$ $-\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{1,i+\frac{1}{2}}^{-}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)+\mathscr{B}_{1,i-\frac{1}{2}}^{+}\left(y_{j}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\right)$
$-\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\mathscr{B}_{2,j+\frac{1}{2}}^{-}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)+\mathscr{B}_{2,j-\frac{1}{2}}^{+}\left(x_{i}^{(\mu)}\right)\boldsymbol{S}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\right).$
We define, for any $1\leq\mu\leq q$,
$\bar{\boldsymbol{U}}_{1,ij}^{(\mu)}:=\frac{1}{\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}}\left[\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}-\boldsymbol{F}_{1}\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\right)\right)\right.+\left.\left(\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}+\boldsymbol{F}_{1}\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\right)\right)\right],$
$\bar{\boldsymbol{U}}_{2,ij}^{(\mu)}:=\frac{1}{\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}}\left[\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}-\boldsymbol{F}_{2}\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\right)\right)\right.+\left.\left(\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}+\boldsymbol{F}_{2}\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\right)\right)\right],$
@@
$\boldsymbol{\Xi}_{2}=\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)\bar{\boldsymbol{U}}_{1,ij}^{(\mu)}+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)\bar{\boldsymbol{U}}_{2,ij}^{(\mu)}.$
Now, we first show $\bar{\rho}_{ij}^{n+1}>0$. Thanks to Lemma 3, we have, for all $1\leq\mu\leq q$, $\bar{\boldsymbol{U}}_{1,ij}^{(\mu)}\in\mathcal{G}_{\rho}$ and $\bar{\boldsymbol{U}}_{2,ij}^{(\mu)}\in\mathcal{G}_{\rho}$. Since $\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}>0$ and $\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}>0$, we obtain $\boldsymbol{\Xi}_{2}\in\mathcal{G}_{\rho}$. According to Lemma 4, we have $\boldsymbol{H}_{1,i+\frac{1}{2}}^{\mu}$, $\boldsymbol{H}_{1,i-\frac{1}{2}}^{\mu}$, $\boldsymbol{H}_{2,j+\frac{1}{2}}^{\mu}$, and $\boldsymbol{H}_{2,j-\frac{1}{2}}^{\mu}$ all belong to $\mathcal{G}_{\rho}$, which means $\boldsymbol{\Xi}_{3}\in\mathcal{G}_{\rho}$. Since the first component of $\boldsymbol{\Xi}_{4}$ is zero, we deduce from (67) that
@@
where we have used the convex decomposition (50) in the above equality, the condition (58) in the second inequality, and the CFL condition (66) in the last inequality.
We then prove that $\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{{\left\|\boldsymbol{B}^{*}\right\|}^{2}}{2}>0$ for any $\boldsymbol{v}^{*},\boldsymbol{B}^{*}\in\mathbb{R}^{3}$. It follows from (67) that
$\bar{\boldsymbol{U}}_{ij}^{n+1}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}=\Pi_{0}+\Pi_{1}+\Pi_{2}+\Pi_{3}+\Pi_{4},$ (68)
@@
$\Pi_{0}:=\bar{\boldsymbol{U}}_{ij}^{n}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2},$
@@
$\Pi_{1}:=$ $\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\mathscr{V}_{1,i+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}\right)\left(\boldsymbol{U}_{i+\frac{1}{2},j}^{-,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right.$
$\left.-\left(\mathscr{V}_{1,i-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)\left(\boldsymbol{U}_{i-\frac{1}{2},j}^{+,\mu}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right]$
@@
$+\frac{\Delta t_{n}}{\Delta y}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left[\left(\mathscr{V}_{2,j+\frac{1}{2}}^{\mu,-}-\hat{\alpha}_{2,j+\frac{1}{2}}^{\mu,-}\right)\left(\boldsymbol{U}_{i,j+\frac{1}{2}}^{\mu,-}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right.$
$\left.-\left(\mathscr{V}_{2,j-\frac{1}{2}}^{\mu,+}+\hat{\alpha}_{2,j-\frac{1}{2}}^{\mu,+}\right)\left(\boldsymbol{U}_{i,j-\frac{1}{2}}^{\mu,+}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)\right],$
@@
$\Pi_{2}:=$ $\frac{\Delta t_{n}}{\Delta x}\sum_{\mu=1}^{q}\omega_{\mu}^{G}\left(\hat{\alpha}_{1,i+\frac{1}{2}}^{\mu,-}+\hat{\alpha}_{1,i-\frac{1}{2}}^{\mu,+}\right)\left(\bar{\boldsymbol{U}}_{1,ij}^{(\mu)}\cdot\boldsymbol{n}^{*}+\frac{\left\|\boldsymbol{B}^{*}\right\|^{2}}{2}\right)$